# AgentXploit: End-to-End Redteaming of Black-Box AI Agents

### v1

@@ Abstract
The strong planning and reasoning capabilities of Large Language Models (LLMs) have fostered the development of agent-based systems capable of leveraging external tools and interacting with increasingly complex environments. However, these powerful features also introduce a critical security risk: indirect prompt injection, a sophisticated attack vector that compromises the core of these agents, the LLM, by manipulating contextual information rather than direct user prompts. In this work, we propose a generic black-box fuzzing framework, AgentXploit, designed to automatically discover and exploit indirect prompt injection vulnerabilities across diverse LLM agents. Our approach starts by constructing a high-quality initial seed corpus, then employs a seed selection algorithm based on Monte Carlo Tree Search (MCTS) to iteratively refine inputs, thereby maximizing the likelihood of uncovering agent weaknesses. We evaluate AgentXploit on two public benchmarks, AgentDojo and VWA-adv, where it achieves 71% and 70% success rates against agents based on o3-mini and GPT-4o, respectively, nearly doubling the performance of baseline attacks. Moreover, AgentXploit exhibits strong transferability across unseen tasks and internal LLMs, as well as promising results against defenses. Beyond benchmark evaluations, we apply our attacks in real-world environments, successfully misleading agents to navigate to arbitrary URLs, including malicious sites.
@@ 1 Introduction
In this work, we propose AgentXploit, the first generic indirect prompt injection assessment method against black-box LLM agents. We draw inspiration from traditional software fuzzing techniques (Miller et al., 1990), which automatically generate test inputs for target software to identify vulnerabilities without requiring access to the software’s internals. We follow the classical fuzzing workflow and design a scalable fuzzing framework for indirect prompt injection attacks on black-box LLM agents. At a high level, given a target LLM agent and a set of seeds for attack instructions, AgentXploit heuristically selects a seed, mutates it, and feeds it to the target agent. Based on the agent’s output, AgentXploit scores the potential and effectiveness of the mutated inputs, adds them to the seed corpus and repeats this process. Fuzzing follows a genetic method that conducts exploration and exploitation in the input space to identify potential vulnerabilities. LLM agents introduce unique challenges to which existing fuzz testing methods cannot be applied: mainly, sparse feedback signals and unique input structure. Under a black-box setting, the only feedback signal available in the LLM agent is whether the target attack has succeeded or not. It is an extremely sparse signal that may downgrade the fuzzing into a random search. To tackle this challenge, we introduce the following three designs: a corpus of high-quality templates, adaptive seed scoring strategies, and a Monte Carlo Tree Search (MCTS)-based seed selection algorithm. The corpus provides initial heuristics, enabling the fuzzing process to have meaningful signals at the early stage. We then introduce an adaptive seed scoring strategy based on attack coverage. It provides intermediate feedback in addition to the final binary success-or-failure feedback, introducing the fuzzing’s exploration effectiveness. Our MCTS-based seed selection algorithm dynamically identifies and prioritizes valuable seeds, improving the exploitation effectiveness. We further design customized mutators for LLM agents’ inputs. As described in  Section 4, the strategies we design are general and can be applied to a variety of proxy and attack tasks.
@@ 2 Related Work
Figure 2: Architecture of AgentXploit and typical process of an indirect prompt injection attack. AgentXploit systematically enhances indirect prompt injection attacks by iteratively refining adversarial prompts. It begins with a high-quality initial corpus of prompt templates, which are tested across various injection tasks to generate initial seeds. Through an iterative fuzzing loop, a Monte Carlo Tree Search (MCTS)-based seed selector identifies promising seeds, a mutator applies transformations, and the modified prompts are evaluated based on attack success and task coverage by the scorer. This adaptive approach ensures scalability and effectiveness across diverse agent architectures and tasks.
@@ 4.1 Overview
Figure 2 illustrates the architecture and the workflow of our proposed framework, AgentXploit. AgentXploit enhances the effectiveness of the indirect prompt attacks by systematically exploring adversarial prompts. The process begins by applying the initial corpus of adversarial prompt templates to the agent across a set of injection tasks, which are combinations of different user tasks and attacker goals, to generate a pool of initial seeds. These seeds then undergo an iterative fuzzing loop. In this loop, a MCTS-based seed selector identifies a promising seed, balancing the dual objectives of exploitation and exploration. Subsequently, a seed mutator randomly selects a mutation method to produce a new variant, which is then tested across the tasks. This variant is subsequently tested across the injection tasks to evaluate its performance. The evaluation involves scoring the new seed based on its success rate in executing attacks and its ability to compromise previously unaffected tasks. Through this adaptive and iterative process, the framework continuously improve the attack, ensuring scalability and effectiveness across a wide range of agents and tasks.
@@ 5 Evaluation
In this section, we comprehensively evaluate the effectiveness of AgentXploit through the following analyses:
@@
    We evaluate AgentXploit on two estabilished agent benchmarks, AgentDojo (Debenedetti et al., 2024) (Section 5.1), representing personal assistant agents, and VWA-adv (Wu et al., 2024b) (Section 5.2), representing web-based agents, covering a variety of agent types and tasks.
@@
    We evaluate the transferability of adversarial prompts generated by AgentXploit across different LLMs and different tasks (Section 5.1&5.2).
@@
    We evaluate the effectiveness of adversarial prompts generated by AgentXploit perform against various defense strategies deployed in the two benchmarks (Section 5.1&5.2).
@@
    We perform an ablation study to understand the contribution of key components of AgentXploit (Section 5.3).
@@ 5.1 Attack Personal Assistant Agents
In this section, we evaluate AgentXploit using the AgentDojo framework (Debenedetti et al., 2024), which is specifically designed for assessing indirect prompt injection attacks and defenses. AgentDojo comprises several components: the environment, which defines an application area for an AI agent along with a set of available tools (such as a workspace environment with email, calendar, and cloud storage access); and the environment state, which tracks data for all applications the agent can interact with. Certain parts of the environment state are specified as placeholders for potential indirect prompt injection attacks. A user task is a natural language user query that the agent is expected to execute within the given environment (e.g., adding an event to a calendar), while an injection task outlines the attacker’s objective (e.g., extracting the user’s credit card information). The collection of user tasks and injection tasks for an specific environment is referred to as a task suite. AgentDojo provides formal evaluation criteria to assess the state of the environment, thereby measuring the success of both user and injection tasks. In our context, a specific attack scenario or an adversarial task is defined as the combination of a user task and an injection task. AgentXploit interacts with AgentDojo by proposing adversarial prompts, which are then inserted into the placeholders in the environment for injection. The agent is subsequently run, and AgentDojo evaluates the success of the user and injection tasks. The success of the injection tasks serves as the attack success signal, providing feedback to AgentXploit.
@@
To evaluate the fuzzing performance and quality of the adversarial prompts generated by AgentXploit, we randomly dividing the adversarial tasks within each suite of AgentDojo into two groups: a fuzzing set and a test set, with 142 and 173 tasks respectively. We utilize GPT-4o-mini as the helper model to mutate the prompts in AgentXploit. We conduct the fuzzing experiment on the fuzzing set for the agent which utilizes the o3-mini model as the backbone due to its state-of-the-art reasoning capabilities. We generate 3 mutated prompts in each iteration and complete a total of 10 fuzzing iterations. Due to the large number of tasks, we randomly sample a quarter of user and injection tasks from each suite to evaluate each newly mutated seeds. For the transferability experiment, we select the 5 seeds with the highest scores. We evaluate the attack performance of the adversarial prompts, against o3-mini, GPT-4o, GPT-4o-mini, and Claude-3.5-Sonnet on the test set. The success rate is computed on the union of the adversarial prompts. According to AgentDojo, the Gemini and DeepSeek families and other open-source models do not fully support the tool call functionality or are not as capable as the aforementioned LLMs. We use the handcrafted adversarial prompts proposed in AgentDojo as the baseline attack. Furthermore, we assess the effectiveness of the generated adversarial prompts against defenses proposed in AgentDojo on the fuzzing set. The defenses includes: pi_detector (ProtectAI, 2024) utilizes a BERT classifier from ProtectAI to detect prompt injection; repeat (lea, 2023) repeats the user instructions after each function call; delimit (Hines et al., 2024) formats all tool outputs with special delimiters and incorporates system prompts to prioritize user instructions. We exclude the tool_filter (Willison, ) defense proposed in AgentDojo due to incompatibility with the o3-mini model. We exclude other defenses due to several key reasons: they struggle to maintain utility, and face issues of high computational costs and adaptability. For example, StruQ (Chen et al., 2024a) is demonstrated only on small open-source models, which lack the capability for agent tasks. Similarly, IsolateGPT (Wu et al., 2025) relies on a system-specific design that cannot be easily adapted to different agent architectures.
@@
Figure 3: Coverage over fuzzing iteration steps achieved by AgentXploit (the solid line) on AgentDojo with two ablation settings (the dashed lines): (1) without the high-quality initial corpus, (2) without the adaptive seed scoring strategy and the MCTS-based seed selection.
Figure 3 presents the coverage progression over the course of the fuzzing iteration steps for AgentXploit. As shown, AgentXploit continuously enhances the performance of the attack, resulting in higher coverage throughout the fuzzing process. In terms of attack success rates, we compare AgentXploit against the baseline handcrafted attacks in AgentDojo, which achieve a success rate of 38%. Our initial high-quality corpus demonstrates a 63% success rate, showcasing its ability to surpass baseline prompts. As fuzzing iterations progress and the adversarial prompts are further refined, AgentXploit achieves a 71% success rate—a significant improvement over both the baseline and the initial corpus. These findings underscore the efficacy of adaptive fuzzing for uncovering injection vulnerabilities in blackbox agents and highlight the effectiveness of targeted search strategies in maximizing the attack performance.
@@
Table 1: The transfer attack success rate of selected adversarial prompts generated by AgentXploit compared with the baseline attacks proposed by AgentDojo (Debenedetti et al., 2024) and VWA-adv (Wu et al., 2024b), against the agents using different backbone LLMs. We run fuzzing against o3-mini on AgentDojo, GPT-4o on VWA-adv.
@@
As shown in Table 1, the success rate results in the first column for o3-mini on the test task set indicate that the generated adversarial prompts transfer effectively across different tasks, even with varying user tasks and injection goals, significantly outperforming the baseline attack—nearly doubling its performance. Comparing performance across rows, we observe that the adversarial prompts transfer well to GPT-4o-mini but perform relatively worse on GPT-4o and Claude-3.5-Sonnet. Furthermore, both the baseline and AgentXploit’s prompts are ineffective against Claude-3.5-Sonnet, as it demonstrates strong robustness in defending against complex adversarial prompts.
@@
Table 2: The attack success rate of selected adversarial prompts generated by AgentXploit on fuzzing task set and o3-mini against four defenses proposed by AgentDojo.
@@
AgentXploit 0.71 0.25 0.12 0.49
The results in Table 2 demonstrate the effectiveness of AgentXploit against defenses compared to the baseline. Checking along the columns, AgentXploit consistently outperforms the baseline, particularly against pi_detector and delimit, indicating that adversarial prompts generated by AgentXploit are more resilient to these defenses. Examining the results along the rows, both the baseline and AgentXploit experience significant drops in success rates when defenses are applied. However, the attacks still maintain high success rates, highlighting the insufficiency of these defenses. Additionally, according to the results, delimit is less effective than pi_detector and repeat, as both AgentXploit and baseline achieve higher success rate against delimit among the defenses.
@@ 5.2 Attacking Web Agents
In this section, we further evaluate AgentXploit on VWA-adv (Wu et al., 2024b). VWA-adv is a set of realistic adversarial tasks based on VisualWebArena (Koh et al., 2024), which serves as a benchmark for evaluating web agents on a set of diverse and complex web-based visual tasks with multi-modal input. Each task in VWA-adv consists of an original task in VisualWebArena and a trigger image or trigger text, which serves as the injection point, along with a targeted adversarial goal as the attacker’s objective. In VWA-adv, attacker goals fall into two categories: illusioning, which misleads agents about object attributes (e.g., changing an object’s color), and goal misdirection, which alters the agent’s intended action (e.g., adding an item to the cart). We focus on the tasks with text trigger. Similar to Section 5.1, we feed the adversarial prompts from AgentXploit to the evaluation framework in VWA-adv, which then returns whether the adversarial task succeeds or not as feedback to AgentXploit.
Similarly, we randomly divide the tasks in VWA-adv into a fuzzing set (99 tasks) and a test set (100 tasks) to evaluate the fuzzing performance and quality of the generated adversarial prompts, respectively. We utilize GPT-4o-mini as the helper model to mutate the prompts in AgentXploit. We run the fuzzing experiment against the agents using GPT-4o on the fuzzing set. We generate 10 mutated prompts per iteration and conduct 10 iterations in total. We use the handcrafted adversarial prompts proposed in VWA-adv as the baseline. We select 5 seeds with the highest scores to conduct the transferability experiment. We evaluate the attack performance of the adversarial prompts against GPT-4o, GPT-4o-mini, Claude-3.5-Sonnet, and Gemini-2-flash-exp on the test set. We further assess the effectiveness against basline defenses proposed in VWA-adv on the fuzzing set. There are three defenses: safety (Hines et al., 2024) utilizes the data delimiter and system prompts to prioritize user instructions; paraphrase (Jain et al., 2023) paraphrases untrusted text to neutralize malicious intent; combined integrates both strategies. While VWA-adv includes one more defense which checks consistency between image and text content, we exclude it since it would substantially increase API calls, making it impractical for real-world use.
@@
Figure 4 shows AgentXploit’s coverage progression during fuzzing. AgentXploit steadily improves attack performance, achieving higher coverage. Compared to baseline attacks in VWA-adv with a success rate of 36%, our high-quality initial corpus starts at 54% and surpasses the baseline. With iterative refinement, AgentXploit reaches 70%, nearly doubling the baseline’s success rate and significantly outperforming both. These results demonstrate AgentXploit’s effectiveness in exposing injection vulnerabilities and optimizing attack performance.
@@
The lower half of Table 1 presents the attack success rates of adversarial prompts from AgentXploit compared to the VWA-adv baseline. The results demonstrate that AgentXploit significantly outperforms the baseline, achieving an absolute success rate improvement of 15% to 40% across different models and tasks except Claude-3.5-Sonnet. This highlights the high quality and effectiveness of the adversarial prompts, as well as their strong transferability. Consistent with Section 5.1, adversarial prompts optimized for GPT do not transfer well to Claude, whereas baseline attacks from VWA-adv achieve higher success rates on Claude compared to other models. Upon manual inspection, we suspect that Claude is more vulnerable to simpler adversarial prompts, differing from the GPT family. Furthermore, the findings reinforce the conclusion from VWA-adv that prompt injection is an effective attack capable of overriding the influence of visual input on the model. It is worth noting that AgentXploit achieve approximately 50% and 60% on GPT-4o-mini and GPT-4o, respectively, suggesting that the instruction hierarchy (Wallace et al., 2024) defense mechanism is not sufficiently effective.
@@
Table 3: The attack success rate of selected adversarial prompts generated by AgentXploit on fuzzing task set and GPT-4o against three defenses proposed by VWA-adv.
@@
AgentXploit 0.60 0.29 0.33 0.27
The evaluation results in Table 3 of AgentXploit highlight a substantial improvement in attack success when no defense mechanisms are applied, achieving a 60% success rate compared to the baseline’s 36%. However, when defenses are introduced, AgentXploit’s performance declines and converges with the baseline. This degradation is likely due to the complexity of the attack prompts, which, while effective in an unprotected setting, struggle against the defenses due to the limited context in VWA-adv. Notably, we observe that the combined defense does not further reduce the attack success rate compared to individual defenses. This suggests that certain attack prompts are inherently more robust and can bypass multiple defenses simultaneously, indicating potential weaknesses in the current defense mechanisms.

### v2

# AgentFuzzer: Generic Black-Box Fuzzing for Indirect Prompt Injection against LLM Agents

@@ Abstract
The strong planning and reasoning capabilities of Large Language Models (LLMs) have fostered the development of agent-based systems capable of leveraging external tools and interacting with increasingly complex environments. However, these powerful features also introduce a critical security risk: indirect prompt injection, a sophisticated attack vector that compromises the core of these agents, the LLM, by manipulating contextual information rather than direct user prompts. In this work, we propose a generic black-box fuzzing framework, AgentFuzzer, designed to automatically discover and exploit indirect prompt injection vulnerabilities across diverse LLM agents. Our approach starts by constructing a high-quality initial seed corpus, then employs a seed selection algorithm based on Monte Carlo Tree Search (MCTS) to iteratively refine inputs, thereby maximizing the likelihood of uncovering agent weaknesses. We evaluate AgentFuzzer on two public benchmarks, AgentDojo and VWA-adv, where it achieves 71% and 70% success rates against agents based on o3-mini and GPT-4o, respectively, nearly doubling the performance of baseline attacks. Moreover, AgentFuzzer exhibits strong transferability across unseen tasks and internal LLMs, as well as promising results against defenses. Beyond benchmark evaluations, we apply our attacks in real-world environments, successfully misleading agents to navigate to arbitrary URLs, including malicious sites.
@@ 1 Introduction
In this work, we propose AgentFuzzer, the first generic indirect prompt injection assessment method against black-box LLM agents. We draw inspiration from traditional software fuzzing techniques (Miller et al., 1990), which automatically generate test inputs for target software to identify vulnerabilities without requiring access to the software’s internals. We follow the classical fuzzing workflow and design a scalable fuzzing framework for indirect prompt injection attacks on black-box LLM agents. At a high level, given a target LLM agent and a set of seeds for attack instructions, AgentFuzzer heuristically selects a seed, mutates it, and feeds it to the target agent. Based on the agent’s output, AgentFuzzer scores the potential and effectiveness of the mutated inputs, adds them to the seed corpus and repeats this process. Fuzzing follows a genetic method that conducts exploration and exploitation in the input space to identify potential vulnerabilities. LLM agents introduce unique challenges to which existing fuzz testing methods cannot be applied: mainly, sparse feedback signals and unique input structure. Under a black-box setting, the only feedback signal available in the LLM agent is whether the target attack has succeeded or not. It is an extremely sparse signal that may downgrade the fuzzing into a random search. To tackle this challenge, we introduce the following three designs: a corpus of high-quality templates, adaptive seed scoring strategies, and a Monte Carlo Tree Search (MCTS)-based seed selection algorithm. The corpus provides initial heuristics, enabling the fuzzing process to have meaningful signals at the early stage. We then introduce an adaptive seed scoring strategy based on attack coverage. It provides intermediate feedback in addition to the final binary success-or-failure feedback, introducing the fuzzing’s exploration effectiveness. Our MCTS-based seed selection algorithm dynamically identifies and prioritizes valuable seeds, improving the exploitation effectiveness. We further design customized mutators for LLM agents’ inputs. As described in  Section 4, the strategies we design are general and can be applied to a variety of proxy and attack tasks.
Differences from GPTFuzzer. GPTFuzzer (Yu et al., 2023) applies fuzzing to jailbreak LLMs via direct prompt injection, it assumes full control over the input and operates in single-turn settings. In contrast, our work targets indirect prompt injection in multi-step agents, where attackers can only influence external content, significantly limiting the capability of the attackers. AgentFuzzer introduces new components, including black-box reward modeling, adaptive seed selection, semantically guided mutators, and carefully designed initial seeds, to address these challenges, making it the first automated black-box framework for attacking LLM-based agents in realistic settings.
@@ 2 Related Work
Figure 2: Architecture of AgentFuzzer and typical process of an indirect prompt injection attack. AgentFuzzer systematically enhances indirect prompt injection attacks by iteratively refining adversarial prompts. It begins with a high-quality initial corpus of prompt templates, which are tested across various injection tasks to generate initial seeds. Through an iterative fuzzing loop, a Monte Carlo Tree Search (MCTS)-based seed selector identifies promising seeds, a mutator applies transformations, and the modified prompts are evaluated based on attack success and task coverage by the scorer. This adaptive approach ensures scalability and effectiveness across diverse agent architectures and tasks.
@@ 4.1 Overview
Figure 2 illustrates the architecture and the workflow of our proposed framework, AgentFuzzer. AgentFuzzer enhances the effectiveness of the indirect prompt attacks by systematically exploring adversarial prompts. The process begins by applying the initial corpus of adversarial prompt templates to the agent across a set of injection tasks, which are combinations of different user tasks and attacker goals, to generate a pool of initial seeds. These seeds then undergo an iterative fuzzing loop. In this loop, a MCTS-based seed selector identifies a promising seed, balancing the dual objectives of exploitation and exploration. Subsequently, a seed mutator randomly selects a mutation method to produce a new variant, which is then tested across the tasks. This variant is subsequently tested across the injection tasks to evaluate its performance. The evaluation involves scoring the new seed based on its success rate in executing attacks and its ability to compromise previously unaffected tasks. Through this adaptive and iterative process, the framework continuously improve the attack, ensuring scalability and effectiveness across a wide range of agents and tasks.
@@ 5 Evaluation
In this section, we comprehensively evaluate the effectiveness of AgentFuzzer through the following analyses:
@@
    We evaluate AgentFuzzer on two estabilished agent benchmarks, AgentDojo (Debenedetti et al., 2024) (Section 5.1), representing personal assistant agents, and VWA-adv (Wu et al., 2024b) (Section 5.2), representing web-based agents, covering a variety of agent types and tasks.
@@
    We evaluate the transferability of adversarial prompts generated by AgentFuzzer across different LLMs and different tasks (Section 5.1&5.2).
@@
    We evaluate the effectiveness of adversarial prompts generated by AgentFuzzer perform against various defense strategies deployed in the two benchmarks (Section 5.1&5.2).
@@
    We perform an ablation study to understand the contribution of key components of AgentFuzzer (Section 5.3).
@@ 5.1 Attack Personal Assistant Agents
In this section, we evaluate AgentFuzzer using the AgentDojo framework (Debenedetti et al., 2024), which is specifically designed for assessing indirect prompt injection attacks and defenses. AgentDojo comprises several components: the environment, which defines an application area for an AI agent along with a set of available tools (such as a workspace environment with email, calendar, and cloud storage access); and the environment state, which tracks data for all applications the agent can interact with. Certain parts of the environment state are specified as placeholders for potential indirect prompt injection attacks. A user task is a natural language user query that the agent is expected to execute within the given environment (e.g., adding an event to a calendar), while an injection task outlines the attacker’s objective (e.g., extracting the user’s credit card information). The collection of user tasks and injection tasks for an specific environment is referred to as a task suite. AgentDojo provides formal evaluation criteria to assess the state of the environment, thereby measuring the success of both user and injection tasks. In our context, a specific attack scenario or an adversarial task is defined as the combination of a user task and an injection task. AgentFuzzer interacts with AgentDojo by proposing adversarial prompts, which are then inserted into the placeholders in the environment for injection. The agent is subsequently run, and AgentDojo evaluates the success of the user and injection tasks. The success of the injection tasks serves as the attack success signal, providing feedback to AgentFuzzer.
@@
To evaluate the fuzzing performance and quality of the adversarial prompts generated by AgentFuzzer, we randomly dividing the adversarial tasks within each suite of AgentDojo into two groups: a fuzzing set and a test set, with 142 and 173 tasks respectively. We utilize GPT-4o-mini as the helper model to mutate the prompts in AgentFuzzer. We conduct the fuzzing experiment on the fuzzing set for the agent which utilizes the o3-mini model as the backbone due to its state-of-the-art reasoning capabilities. We generate 3 mutated prompts in each iteration and complete a total of 10 fuzzing iterations. Due to the large number of tasks, we randomly sample a quarter of user and injection tasks from each suite to evaluate each newly mutated seeds. For the transferability experiment, we select the 5 seeds with the highest scores. We evaluate the attack performance of the adversarial prompts, against o3-mini, GPT-4o, GPT-4o-mini, and Claude-3.5-Sonnet on the test set. The success rate is computed on the union of the adversarial prompts. According to AgentDojo, the Gemini and DeepSeek families and other open-source models do not fully support the tool call functionality or are not as capable as the aforementioned LLMs. We use the handcrafted adversarial prompts proposed in AgentDojo as the baseline attack. Furthermore, we assess the effectiveness of the generated adversarial prompts against defenses proposed in AgentDojo on the fuzzing set. The defenses includes: pi_detector (ProtectAI, 2024) utilizes a BERT classifier from ProtectAI to detect prompt injection; repeat (lea, 2023) repeats the user instructions after each function call; delimit (Hines et al., 2024) formats all tool outputs with special delimiters and incorporates system prompts to prioritize user instructions. We exclude the tool_filter (Willison, ) defense proposed in AgentDojo due to incompatibility with the o3-mini model. We exclude other defenses due to several key reasons: they struggle to maintain utility, and face issues of high computational costs and adaptability. For example, StruQ (Chen et al., 2024a) is demonstrated only on small open-source models, which lack the capability for agent tasks. Similarly, IsolateGPT (Wu et al., 2025) relies on a system-specific design that cannot be easily adapted to different agent architectures.
@@
Figure 3: Coverage over fuzzing iteration steps achieved by AgentFuzzer (the solid line) on AgentDojo with two ablation settings (the dashed lines): (1) without the high-quality initial corpus, (2) without the adaptive seed scoring strategy and the MCTS-based seed selection.
Figure 3 presents the coverage progression over the course of the fuzzing iteration steps for AgentFuzzer. As shown, AgentFuzzer continuously enhances the performance of the attack, resulting in higher coverage throughout the fuzzing process. In terms of attack success rates, we compare AgentFuzzer against the baseline handcrafted attacks in AgentDojo, which achieve a success rate of 38%. Our initial high-quality corpus demonstrates a 63% success rate, showcasing its ability to surpass baseline prompts. As fuzzing iterations progress and the adversarial prompts are further refined, AgentFuzzer achieves a 71% success rate—a significant improvement over both the baseline and the initial corpus. These findings underscore the efficacy of adaptive fuzzing for uncovering injection vulnerabilities in blackbox agents and highlight the effectiveness of targeted search strategies in maximizing the attack performance.
@@
Table 1: The transfer attack success rate of selected adversarial prompts generated by AgentFuzzer compared with the baseline attacks proposed by AgentDojo (Debenedetti et al., 2024) and VWA-adv (Wu et al., 2024b), against the agents using different backbone LLMs. We run fuzzing against o3-mini on AgentDojo, GPT-4o on VWA-adv.
@@
As shown in Table 1, the success rate results in the first column for o3-mini on the test task set indicate that the generated adversarial prompts transfer effectively across different tasks, even with varying user tasks and injection goals, significantly outperforming the baseline attack—nearly doubling its performance. Comparing performance across rows, we observe that the adversarial prompts transfer well to GPT-4o-mini but perform relatively worse on GPT-4o and Claude-3.5-Sonnet. Furthermore, both the baseline and AgentFuzzer’s prompts are ineffective against Claude-3.5-Sonnet, as it demonstrates strong robustness in defending against complex adversarial prompts.
@@
Table 2: The attack success rate of selected adversarial prompts generated by AgentFuzzer on fuzzing task set and o3-mini against four defenses proposed by AgentDojo.
@@
AgentFuzzer 0.71 0.25 0.12 0.49
The results in Table 2 demonstrate the effectiveness of AgentFuzzer against defenses compared to the baseline. Checking along the columns, AgentFuzzer consistently outperforms the baseline, particularly against pi_detector and delimit, indicating that adversarial prompts generated by AgentFuzzer are more resilient to these defenses. Examining the results along the rows, both the baseline and AgentFuzzer experience significant drops in success rates when defenses are applied. However, the attacks still maintain high success rates, highlighting the insufficiency of these defenses. Additionally, according to the results, delimit is less effective than pi_detector and repeat, as both AgentFuzzer and baseline achieve higher success rate against delimit among the defenses.
@@ 5.2 Attacking Web Agents
In this section, we further evaluate AgentFuzzer on VWA-adv (Wu et al., 2024b). VWA-adv is a set of realistic adversarial tasks based on VisualWebArena (Koh et al., 2024), which serves as a benchmark for evaluating web agents on a set of diverse and complex web-based visual tasks with multi-modal input. Each task in VWA-adv consists of an original task in VisualWebArena and a trigger image or trigger text, which serves as the injection point, along with a targeted adversarial goal as the attacker’s objective. In VWA-adv, attacker goals fall into two categories: illusioning, which misleads agents about object attributes (e.g., changing an object’s color), and goal misdirection, which alters the agent’s intended action (e.g., adding an item to the cart). We focus on the tasks with text trigger. Similar to Section 5.1, we feed the adversarial prompts from AgentFuzzer to the evaluation framework in VWA-adv, which then returns whether the adversarial task succeeds or not as feedback to AgentFuzzer.
Similarly, we randomly divide the tasks in VWA-adv into a fuzzing set (99 tasks) and a test set (100 tasks) to evaluate the fuzzing performance and quality of the generated adversarial prompts, respectively. We utilize GPT-4o-mini as the helper model to mutate the prompts in AgentFuzzer. We run the fuzzing experiment against the agents using GPT-4o on the fuzzing set. We generate 10 mutated prompts per iteration and conduct 10 iterations in total. We use the handcrafted adversarial prompts proposed in VWA-adv as the baseline. We select 5 seeds with the highest scores to conduct the transferability experiment. We evaluate the attack performance of the adversarial prompts against GPT-4o, GPT-4o-mini, Claude-3.5-Sonnet, and Gemini-2-flash-exp on the test set. We further assess the effectiveness against basline defenses proposed in VWA-adv on the fuzzing set. There are three defenses: safety (Hines et al., 2024) utilizes the data delimiter and system prompts to prioritize user instructions; paraphrase (Jain et al., 2023) paraphrases untrusted text to neutralize malicious intent; combined integrates both strategies. While VWA-adv includes one more defense which checks consistency between image and text content, we exclude it since it would substantially increase API calls, making it impractical for real-world use.
@@
Figure 4 shows AgentFuzzer’s coverage progression during fuzzing. AgentFuzzer steadily improves attack performance, achieving higher coverage. Compared to baseline attacks in VWA-adv with a success rate of 36%, our high-quality initial corpus starts at 54% and surpasses the baseline. With iterative refinement, AgentFuzzer reaches 70%, nearly doubling the baseline’s success rate and significantly outperforming both. These results demonstrate AgentFuzzer’s effectiveness in exposing injection vulnerabilities and optimizing attack performance.
@@
The lower half of Table 1 presents the attack success rates of adversarial prompts from AgentFuzzer compared to the VWA-adv baseline. The results demonstrate that AgentFuzzer significantly outperforms the baseline, achieving an absolute success rate improvement of 15% to 40% across different models and tasks except Claude-3.5-Sonnet. This highlights the high quality and effectiveness of the adversarial prompts, as well as their strong transferability. Consistent with Section 5.1, adversarial prompts optimized for GPT do not transfer well to Claude, whereas baseline attacks from VWA-adv achieve higher success rates on Claude compared to other models. Upon manual inspection, we suspect that Claude is more vulnerable to simpler adversarial prompts, differing from the GPT family. Furthermore, the findings reinforce the conclusion from VWA-adv that prompt injection is an effective attack capable of overriding the influence of visual input on the model. It is worth noting that AgentFuzzer achieve approximately 50% and 60% on GPT-4o-mini and GPT-4o, respectively, suggesting that the instruction hierarchy (Wallace et al., 2024) defense mechanism is not sufficiently effective.
@@
Table 3: The attack success rate of selected adversarial prompts generated by AgentFuzzer on fuzzing task set and GPT-4o against three defenses proposed by VWA-adv.
@@
AgentFuzzer 0.60 0.29 0.33 0.27
The evaluation results in Table 3 of AgentFuzzer highlight a substantial improvement in attack success when no defense mechanisms are applied, achieving a 60% success rate compared to the baseline’s 36%. However, when defenses are introduced, AgentFuzzer’s performance declines and converges with the baseline. This degradation is likely due to the complexity of the attack prompts, which, while effective in an unprotected setting, struggle against the defenses due to the limited context in VWA-adv. Notably, we observe that the combined defense does not further reduce the attack success rate compared to individual defenses. This suggests that certain attack prompts are inherently more robust and can bypass multiple defenses simultaneously, indicating potential weaknesses in the current defense mechanisms.